2.   The data set below is a sample of the Mathematics test scores of 10 students:

56, 96, 78, 67, 60, 69, 85, 90, 89, 72          

(a) Find the mean and median of the given test scores.

(b) Is there a mode value for these scores?   Why or why not?

(c) Find the range and standard deviation (nearest hundredth) of these scores.

(d) Find the percentile rank of 78

(e) What percent of these scores are within 1 standard deviation from the mean?

2.   The data set below is a sample of the Mathematics test scores of 10 students:

56, 96, 78, 67, 60, 69, 85, 90, 89, 72          

(a) Find the mean and median of the given test scores.

(b) Is there a mode value for these scores?   Why or why not?

(c) Find the range and standard deviation (nearest hundredth) of these scores.

(d) Find the percentile rank of 78

(e) What percent of these scores are within 1 standard deviation from the mean?
Answer #1

2) ascending order values given below

56,60,67,69,72,78,85,89,90,96

Mean = sum of all values/ number of values

= 762/ 10

= 76.2

Median = 5th value + 6th value / 2

= 72+78/2

= 75

B) Mode is a value of maximum repeated value.here,no value repeat more than once.so,there is no mode

C) range = maximum - minimum

= 96 -56

= 40